A charges raindrop carrying 1.0 X 10^-5 C experience an electric force of 0.30N in +X direction. What is the electric field location. (b) What would be the force on a -5.0 X 10^-5 C drop at the same location.
Answer #1

Here ,

charge ,q = 1 *10^-5 C

force , F = 0.30 i N

a) let the electric field is E

E * q = electric force

1 *10^-5 * E = 0.30 i

E = 3 *10^4 i N/C

the electric field is 3 *10^4 N/C in the positive x - direction

b)

force = charge * electric field

force = -5 *10^-5 * 3 *10^4 i

force = -1.5 i N

the force on this charge is 1.5 N in the negative x - direction